OpenStudy (anonymous):

Estelle is raking leaves in her yard. After 1 hour, she has 2 bags of leaves. After 3 hours, she has 6 bags of leaves. Which graph represents Estelle's leaf raking progress? How would this graph look?

OpenStudy (anonymous):

Please help and explain :(

OpenStudy (anonymous):

@mathstudent55 @mathmath333 @iambatman

OpenStudy (mathmath333):

if u observe the hours and bags of leaves ratio is h/b=1/2=3/6.....=n/2n therefore take coordinate (x,y)=(n,2n), u can take n as n=1,2,3,4,5,6........ and plot the graph it will look like this